# Anomalous and linear holographic hard wall models for light unflavored mesons

### v1

@@ Abstract
In this work we consider anomalous and linear holographic hard wall (HW) models for light unflavored mesons inspired by the AdS/CFT correspondence. The anomalous dimensions depend on the logarithm of the spin $S$ of the meson state and come from a semiclassical analysis of gauge/string duality. The anomalous HW model produces very good masses and good Regge trajectories for mesons compared with PDG data. Inspired by this anomalous HW model we also propose a phenomenological modification of the dimension of the boundary operators such that the model produces asymptotic linear Regge trajectories.
@@ I Introduction
The AdS/CFT correspondence Aharony:1999ti establishes relations between Yang-Mills (YM) and higher dimensional gravitational theories. One of the amazing properties of the correspondence is that if YM theories are strongly coupled, the gravitational one is weak coupled allowing for the solution of a myriad of phenomena from the quark-gluon plasma to condensed matter (for reviews, see e.g., Casalderrey-Solana:2011dxg ; Zaanen:2015oix ). In particular, hadronic spectra were described by different proposals as the holographic hard wall (HW) and the soft wall (SW) models Boschi-Filho:2002xih ; Boschi-Filho:2002wdj ; deTeramond:2005su ; Erlich:2005qh ; Boschi-Filho:2005xct ; Karch:2006pv ; Colangelo:2008us . The HW model Boschi-Filho:2002xih ; Boschi-Filho:2002wdj ; deTeramond:2005su ; Erlich:2005qh ; Boschi-Filho:2005xct is well known for producing reasonable hadronic masses but non-linear (parabolic) Regge trajectories. On the other side, the SW model Karch:2006pv ; Colangelo:2008us also give good masses with the advantage of generating linear Regge trajectories.
@@
In this work, we are going to consider the inclusion of anomalous dimensions for the boundary operators describing light unflavored mesons with spins $S\geq 1$. These anomalous dimensions depend on the logarithm of the spin $S$ of the mesonic state inspired by a semiclasical analysis of gauge/string dualities Gubser:2002tv and was used recently to discuss the case of even spins glueballs Costa-Silva:2023vuu . We find very good masses and Regge trajectories for the studied mesons when compared with PDG data PDG:2022pth , better than the ones from the HW and SW models. Furthermore, we discuss a phenomenological modification of the dimension of the boundary operators representing the mesonic states in order to obtain asymptotic linear Regge trajectories, similar to the case of glueballs. Actually, the linearization of the mesonic states require an additional parameter to obtain the best fit to PDG data. For a recent account on non-holographic meson Regge trajectories, see e.g. Chen:2021kfw .
@@
This work is organized as follows. In Section II, we review the HW model for light unflavored mesons and in Section III we discuss the introduction of anomalous dimensions in the HW model and obtain the spectra for the mesonic families starting with the states of the $\rho(770)$, $\omega(782)$, $b_{1}(1235)$, and $h_{1}(1170)$. Then, in Section IV we modify the dimension of the mesonic boundary operators such that we find asymptotic linear Regge trajectories for the mesons already discussed in the previous section. Finally, in Section V we present our conclusions.
@@ II Original HW model for mesons
In this work we are interested in studying the spectra of light unflavored mesons, and their Regge trajectories using anomalous and linear holographic HW models. These mesons can be classified in four families according to their behavior under $\mathcal{CP}$ transformations (even or odd), isospin $I=0$ and $1$, and spins $S\geq 1$. These four families imply four distinct Regge trajectories.
In order to describe such objects using the holographic HW model, we need to study a massive vector field in $AdS_{5}$. We start with the action
@@
where $F_{mn}=\partial_{m}A_{n}-\partial_{n}A_{m}$ is the stress tensor of the Proca field with mass $M_{5}$ in the five-dimensional AdS space. Moreover, $g$ is the determinant of the metric of $AdS_{5}$,
@@
This is Bessel equation, whose solutions are linear combination of Bessel ($J_{\nu}$) and Neumann ($Y_{\nu})$ functions of order $\nu=\sqrt{1+(M_{5}R)^{2}}$. Since we are interested only in regular solutions in the bulk, we will consider just the Bessel solution
@@
On the other hand, the mesons with spin $S$ with $S=1,2,3,...$, that we are interested in are described by operators in the boundary Minkowski space as Gubser:2002tv  [^1] It is also possible to alternatively write the mesons with spin $S$ as deTeramond:2005su ${\cal O}^{\mu\,\mu_{1}...\mu_{S}}_{2+S}=\;\bar{\psi}\gamma^{\mu}D_{\{\mu_{1}\dots}D_{\mu_{S}\}}\psi\;;\qquad(S=1,2,3,...)\,,$ (8) such that $\Delta=2+S.$ (9) However, in this case, the results for the AHW and LAHW are not so good as the ones discussed in Sections III and IV, with $\Delta=3+S$.
${\cal O}^{\mu_{1}...\mu_{S}}_{3+S}=\;\bar{\psi}D_{\{\mu_{1}\dots}D_{\mu_{S}\}}\psi\;;\qquad(S=1,2,3,...)\,,$ (10)
where $\psi$ and $\bar{\psi}$ represent quark and antiquark fields while $D_{\mu_{i}}$ with $i=1,\dots,S$ are symmetrized covariant derivatives. Then, the canonical dimensions of these operators are given by
@@
The link between the operators $\cal O$ and string excitations at $AdS_{5}$ can be constructed from its conformal dimension. From the AdS/CFT correspondence, the vector field in $AdS_{5}$ that we are describing is a $1$-form, so its holographic mass $M_{5}$ is related to the conformal dimension $\Delta$ of boundary operators $\cal O$, is given by
@@
The HW model consists in the holographic relation between fields in AdS space with an IR cutoff that breaks the conformal symmetry of the boundary theory. That means a hard cutoff in the holographic $z$ coordinate, so we are working in an AdS slice, given by $0\leq z\leq z_{\rm max}$, where
$z_{\rm max}=\frac{1}{\Lambda_{\rm QCD}},$ (15)
with $\Lambda_{\rm QCD}$ a convenient energy scale related to be fixed later. Then the solution of the vector bulk field, Eq. (7), assumes the form
@@
where $m_{\nu,k}$ is identified with the mass of the boundary operator $\cal O$ that lives in the four-dimensional Minkowski space, describing mesons. Note that $k=1,2,3...$ denote radial excitations, where $k=1$ is the corresponding ground state. In this work we are interested in angular momenta excitations, so we take $k=1$ for all the objects from now on.
@@
$J_{\nu}\bigg{(}\frac{m_{\nu,1}}{\Lambda_{\rm QCD}}\bigg{)}=0.$ (17)
@@
$m_{\nu,1}=\chi_{\nu,1}\Lambda_{\rm QCD};\qquad\qquad(\nu=S+1).$ (18)
@@
taking $m_{2,1}$ as an input from PDG for the states $\rho(770)$, $\omega(782)$, $b_{1}(1235)$, and $h_{1}(1170)$. The corresponding values of $\Lambda_{\rm QCD}$ found here for each mesonic family are shown in Table 2. The meson masses obtained from the HW model, using Eq. (18), are shown in Table 1, together with the SW model predictions with the relative deviations with respect to PDG data PDG:2022pth . In Table 2, we also show the equations for the Regge trajectories found from PDG data for the same mesonic families.
@@
The process described above is known as the holographic HW model and was used in the past to obtain hadronic spectra among other properties. It is a well-known fact that this model implies non-linear Regge trajectories.
@@
In the next section we develop an improvement of the HW model, by considering anomalous dimensions, in order to obtain better masses compared with the PDG ones and better Regge trajectories for these meson families.
Table 1: Masses in MeV for light unflavored mesons classified under even or odd $\mathcal{CP}$ and with isospin $I=0,1$, and spins $S\geq 1$ from PDG PDG:2022pth , from the HW Boschi-Filho:2002xih ; Boschi-Filho:2002wdj ; deTeramond:2005su ; Erlich:2005qh ; Boschi-Filho:2005xct and SW Karch:2006pv ; Colangelo:2008us models, and the percentual deviations $\delta_{\rm HW}$ and $\delta_{\rm SW}$ of the masses from these two models compared with PDG data. The HW masses are determined from Eq. (18), using the values of $\Lambda_{\rm QCD}$ for each family (same $\mathcal{CP}$ parity and Isospin) shown in Table 2, and the masses from the SW model come from the relation $m_{S}^{2}\sim S$ Karch:2006pv . The mass of the state with $J=S=1$ for each family is used as an input for the HW and SW models.
$J^{PC}$ ${CP}$ $I$ Meson PDG HW $\delta_{\rm HW}$ SW $\delta_{\rm SW}$
$1^{--}$ even 1 $\rho(770)$ $775.26\pm 0.25$ $775.26$ $0\%$ $775.26$ $0\%$
$2^{++}$ even 1 $a_{2}(1320)$ $1316.9\pm 0.9$ $963.13$ $25.9\%$ $1096.38$ $16.7\%$
$3^{--}$ even 1 $\rho_{3}(1690)$ $1688.8\pm 2.1$ $1145.52$ $32.2\%$ $1342.79$ $20.5\%$
$4^{++}$ even 1 $a_{4}(1970)$ $1967\pm 16$ $1324.12$ $32.7\%$ $1550.52$ $21.2\%$
$5^{--}$ even 1 $\rho_{5}(2350)$ $2330\pm 35$ $1499.93$ $35.6\%$ $1733.53$ $25.6\%$
$6^{++}$ even 1 $a_{6}(2450)$ $2450\pm 130$ $1673.57$ $31.7\%$ $1898.99$ $22.5\%$
$1^{--}$ even 0 $\omega(782)$ $782.66\pm 0.13$ $782.66$ $0\%$ $782.66$ $0\%$
$2^{++}$ even 0 $f_{2}(1270)$ $1275.5\pm 0.8$ $972.31$ $23.8\%$ $1106.85$ $13.2\%$
$3^{--}$ even 0 $\omega_{3}(1670)$ $1667\pm 4$ $1156.44$ $30.6\%$ $1355.61$ $18.7\%$
$4^{++}$ even 0 $f_{4}(2050)$ $2018\pm 11$ $1336.74$ $33.8\%$ $1565.32$ $22.4\%$
$5^{--}$ even 0 $\omega_{5}(2250)$ $2250\pm 70$ $1514.23$ $32.7\%$ $1750.08$ $22.2\%$
$6^{++}$ even 0 $f_{6}(2510)$ $2465\pm 50$ $1689.52$ $31.5\%$ $1917.12$ $22.2\%$
$1^{+-}$ odd 1 $b_{1}(1235)$ $1229.5\pm 3.2$ $1229.5$ $0\%$ $1229.5$ $0\%$
$2^{-+}$ odd 1 $\pi_{2}(1670)$ $1670.6\pm 2.05$ $1527.41$ $8.57\%$ $1738.78$ $4.08\%$
$3^{+-}$ odd 1 $b_{3}(2030)$ $2032\pm 12$ $1816.70$ $10.6\%$ $2129.56$ $4.80\%$
$4^{-+}$ odd 1 $\pi_{4}(2250)$ $2250\pm 15$ $2099.95$ $6.67\%$ $2459.00$ $9.29\%$
$1^{+-}$ odd 0 $h_{1}(1170)$ $1166\pm 6$ $1166$ $0\%$ $1166$ $0\%$
$2^{-+}$ odd 0 $\eta_{2}(1645)$ $1617\pm 5$ $1352$ $16.4\%$ $1648.97$ $1.98\%$
$3^{+-}$ odd 0 $h_{3}(2025)$ $2025\pm 20$ $1709$ $15.6\%$ $2019.57$ $0.27\%$
$4^{-+}$ odd 0 $\eta_{4}(2330)$ $2328\pm 38$ $1882$ $19.2\%$ $2332.00$ $0.17\%$
Table 2: The Regge trajectory in GeV units for each light unflavored meson family (same $\mathcal{CP}$ parity and same isospin $I$) from the best fit of PDG data shown in Table 1. We also show the mass of the state with $J=S=1$ for each family, and the corresponding $\Lambda_{\rm QCD}$ value, both in MeV, determined from Eq. (19).
$J^{PC}$ ${CP}$ $I$ Meson PDG (MeV) PDG Regge Trajectory (GeV) $\Lambda_{\rm QCD}$ (MeV)
$1^{--}$ even 1 $\rho(770)$ $775.26\pm 0.25$ $J=(0.88\pm 0.04)M^{2}+0.47\pm 0.15.$ $150.96$
$1^{--}$ even 0 $\omega(782)$ $782.66\pm 0.13$ $J=(0.90\pm 0.02)M^{2}+0.47\pm 0.07$ 152.24
$1^{+-}$ odd 1 $b_{1}(1235)$ $1229.5\pm 3.2$ $J=(0.83\pm 0.04)M^{2}-0.30\pm 0.16$ 239.41
$1^{+-}$ odd 0 $h_{1}(1170)$ $1166\pm 6$ $J=(0.73\pm 0.02)M^{2}+0.03\pm 0.06$ 227.04

### v2

@@ Abstract
[…]
@@ I Introduction
The AdS/CFT correspondence Aharony:1999ti establishes relations between Yang-Mills (YM) and higher dimensional gravitational theories. One of the amazing properties of this correspondence is that if the YM theories are strongly coupled, the dual gravitational one is weak coupled allowing for the solution of a myriad of phenomena from the quark-gluon plasma to condensed matter (for reviews, see e.g., Casalderrey-Solana:2011dxg ; Zaanen:2015oix ). In particular, hadronic spectra were described by different proposals such as the holographic hard wall (HW) and the soft wall (SW) models Polchinski:2001tt ; Polchinski:2002jw ; Boschi-Filho:2002xih ; Boschi-Filho:2002wdj ; deTeramond:2005su ; Erlich:2005qh ; Boschi-Filho:2005xct ; Karch:2006pv ; Colangelo:2008us . The HW model Boschi-Filho:2002xih ; Boschi-Filho:2002wdj ; deTeramond:2005su ; Erlich:2005qh ; Boschi-Filho:2005xct is well known for producing reasonable hadronic masses but non-linear (parabolic) Regge trajectories. On the other side, the SW model Karch:2006pv ; Colangelo:2008us also give good masses with the advantage of generating linear Regge trajectories. For other interesting QCD properties described by the HW model, see e.g. Boschi-Filho:2005nmp ; Boschi-Filho:2006hfm ; Herzog:2006ra ; BallonBayona:2007vp ; Grigoryan:2007my ; Kwee:2007dd ; Grigoryan:2007wn ; Jo:2009xr ; Costa:2013uia ; Craps:2013iaa ; Mamo:2015dea ; Ballon-Bayona:2017bwk .
@@
In this work, we are going to consider the inclusion of anomalous dimensions for the boundary operators describing light unflavored mesons with spins $S\geq 1$ in the HW model. These anomalous dimensions depend on the logarithm of the spin $S$ of the mesonic state inspired by a semiclasical analysis of gauge/string dualities Gubser:2002tv and was used recently to discuss the case of even spin glueballs Costa-Silva:2023vuu . We find very good masses and Regge trajectories for the studied mesons when compared with PDG data PDG:2022pth , better than the ones from the HW and SW models. Furthermore, we discuss a phenomenological modification of the dimension of the boundary operators representing the mesonic states in order to obtain asymptotically linear Regge trajectories, similar to the case of glueballs. Actually, the linearization of the mesonic states require an additional parameter to obtain the best fit to PDG data. For a recent account on non-holographic meson Regge trajectories, see e.g. Chen:2021kfw . For later reference, we present in Table 1 the masses of the meson families starting with the states $\rho(770)$, $\omega(782)$, $b_{1}(1235)$, and $h_{1}(1170)$, with their corresponding Regge trajectories. These families are classified according to their $PC$ parity (even when $P$ and $C$ have the same sign or odd with opposite signs) and isospins $I=0,1$. Note that, according to PDG notation, mesons are classified as $J^{PC}$, where $J$ is the total angular momenta, $P$ is the parity and $C$ the charge conjugation of the state. In our notation, we simply call the total angular momenta as the “spin” $S$, so that $J=S$. The $G$ parity ($G=(-)^{I+S}$) of these mesons is also included in this table following PDG notation but is not relevant in the present analysis.
$I^{G}(J^{PC})$ Meson PDG (MeV) PDG Regge trajetories (GeV)
$1^{+}(1^{--})$ $\rho(770)$ $775.26\pm 0.25$
$1^{-}(2^{++})$ $a_{2}(1320)$ $1316.9\pm 0.9$
$1^{+}(3^{--})$ $\rho_{3}(1690)$ $1688.8\pm 2.1$ $J=(0.88\pm 0.04)M^{2}+0.47\pm 0.15$
$1^{-}(4^{++})$ $a_{4}(1970)$ $1967\pm 16$
$1^{+}(5^{--})$ $\rho_{5}(2350)$ $2330\pm 35$
$1^{-}(6^{++})$ $a_{6}(2450)$ $2450\pm 130$
$0^{-}(1^{--})$ $\omega(782)$ $782.66\pm 0.13$
$0^{+}(2^{++})$ $f_{2}(1270)$ $1275.5\pm 0.8$
$0^{-}(3^{--})$ $\omega_{3}(1670)$ $1667\pm 4$ $J=(0.90\pm 0.02)M^{2}+0.47\pm 0.07$
$0^{+}(4^{++})$ $f_{4}(2050)$ $2018\pm 11$
$0^{-}(5^{--})$ $\omega_{5}(2250)$ $2250\pm 70$
$0^{+}(6^{++})$ $f_{6}(2510)$ $2465\pm 50$
$1^{+}(1^{+-})$ $b_{1}(1235)$ $1229.5\pm 3.2$
$1^{-}(2^{-+})$ $\pi_{2}(1670)$ $1670.6\pm 2.05$ $J=(0.83\pm 0.04)M^{2}-0.30\pm 0.16$
$1^{+}(3^{+-})$ $b_{3}(2030)$ $2032\pm 12$
$1^{-}(4^{-+})$ $\pi_{4}(2250)$ $2250\pm 15$
$0^{-}(1^{+-})$ $h_{1}(1170)$ $1166\pm 6$
$0^{+}(2^{-+})$ $\eta_{2}(1645)$ $1617\pm 5$ $J=(0.73\pm 0.02)M^{2}+0.03\pm 0.06$
$0^{-}(3^{+-})$ $h_{3}(2025)$ $2025\pm 20$
$0^{+}(4^{-+})$ $\eta_{4}(2330)$ $2328\pm 38$
Table 1: Masses in MeV for light unflavored mesons classified under even or odd $PC$, with isospins $I=0,1$, and spins $J=S\geq 1$ from PDG PDG:2022pth , and their corresponding Regge trajectories in GeV.
@@
This work is organized as follows. In Section II, we review the HW model for light unflavored mesons and in Section III we discuss the introduction of anomalous dimensions in the HW model and obtain the spectra for the mesonic families starting with the states of the $\rho(770)$, $\omega(782)$, $b_{1}(1235)$, and $h_{1}(1170)$. Then, in Section IV we modify the dimension of the mesonic boundary operators such that we find asymptotically linear Regge trajectories for the mesons already discussed in the previous section. Finally, in Section V we present our conclusions.
@@ II Original HW model for mesons
In this work we are interested in studying the spectra of light unflavored mesons, and their Regge trajectories using anomalous and linear holographic HW models. These mesons can be classified in four families according to their behavior under ${PC}$ transformations (even or odd), isospin $I=0$ or $1$, and spins $S\geq 1$. These four families imply four distinct Regge trajectories.
In order to describe such objects using the holographic HW model, we need to study a massive vector field in $AdS_{5}$. We start with the action Aharony:1999ti
@@
where $F_{mn}=\partial_{m}A_{n}-\partial_{n}A_{m}$ is the stress tensor of the Proca field $A_{m}$ with mass $M_{5}$ in the five-dimensional AdS space. Moreover, $g$ is the determinant of the metric of $AdS_{5}$,
@@
This is Bessel equation, whose solutions are linear combinations of Bessel ($J_{\nu}$) and Neumann ($Y_{\nu})$ functions both of order $\nu=\sqrt{1+(M_{5}R)^{2}}$. Since we are interested only in regular solutions in the bulk, we will consider just the Bessel solution
@@
On the other hand, the mesons with spin $S=1,2,3,...$, that we are interested in are described by operators in the boundary Minkowski space as Gubser:2002tv
${\cal O}_{\mu_{1}...\mu_{S}}^{3+S}=\;\bar{\psi}D_{\{\mu_{1}\dots}D_{\mu_{S}\}}\psi\;;\qquad(S=1,2,3,...)\,,$ (8)
where $\psi$ and $\bar{\psi}$ represent quark and antiquark fields while $D_{\mu_{i}}$ with $i=1,\dots,S$, are symmetrized covariant derivatives Operador . Then, the canonical dimensions of these operators are given by
@@
The link between the operators $\cal O$ and string excitations in $AdS_{5}$ can be constructed from its conformal dimension. From the AdS/CFT correspondence, the vector field in $AdS_{5}$ that we are describing is a $1$-form, so its holographic mass $M_{5}$ is related to the conformal dimension $\Delta$ of boundary operators $\cal O$, and given by Aharony:1999ti
@@
The HW model consists in the holographic relation between bulk fields in AdS space with an infrared (IR) cutoff that breaks the conformal symmetry of the boundary theory. That means a hard cutoff in the holographic $z$ coordinate, so we are working in an AdS slice, given by $0\leq z\leq z_{\rm max}$, where
$z_{\rm max}=\frac{1}{\Lambda_{\rm HW}},$ (13)
with $\Lambda_{\rm HW}$ a convenient energy scale in the HW model to be fixed later (see Eq. (17) below). Then, the solution of the vector bulk field, Eq. (7), assumes the form
@@
where $m_{\nu,k}$ is identified with the mass of the boundary operator $\cal O$ that lives in the four-dimensional Minkowski space, describing mesons. Note that $k=1,2,3...,$ denote radial excitations, where $k=1$ is the corresponding ground state. In this work we are interested in angular momenta excitations, so we take $k=1$ for all the objects from now on.
$I^{G}(J^{PC})$ Meson HW $\delta_{\rm HW}$ $\Lambda_{\rm HW}$ HW Regge Trajectories (GeV) SW $\delta_{\rm SW}$
$1^{+}(1^{--})$ $\rho(770)$ $775.26$ $0\%$ $775.26$ $0\%$
$1^{-}(2^{++})$ $a_{2}(1320)$ $963.13$ $25.9\%$ $1096.38$ $16.7\%$
$1^{+}(3^{--})$ $\rho_{3}(1690)$ $1145.52$ $32.2\%$ 150.96 $J=(2.25\pm 0.10)M^{2}-0.12\pm 0.19$ $1342.79$ $20.5\%$
$1^{-}(4^{++})$ $a_{4}(1970)$ $1324.12$ $32.7\%$ $1550.52$ $21.2\%$
$1^{+}(5^{--})$ $\rho_{5}(2350)$ $1499.93$ $35.6\%$ $1733.53$ $25.6\%$
$1^{-}(6^{++})$ $a_{6}(2450)$ $1673.57$ $31.7\%$ $1898.99$ $22.5\%$
$0^{-}(1^{--})$ $\omega(782)$ $782.66$ $0\%$ $782.66$ $0\%$
$0^{+}(2^{++})$ $f_{2}(1270)$ $972.31$ $23.8\%$ $1106.85$ $13.2\%$
$0^{-}(3^{--})$ $\omega_{3}(1670)$ $1156.44$ $30.6\%$ 152.24 $J=(2.21\pm 0.10)M^{2}-0.12\pm 0.19$ $1355.61$ $18.7\%$
$0^{+}(4^{++})$ $f_{4}(2050)$ $1336.74$ $33.8\%$ $1565.32$ $22.4\%$
$0^{-}(5^{--})$ $\omega_{5}(2250)$ $1514.23$ $32.7\%$ $1750.08$ $22.2\%$
$0^{+}(6^{++})$ $f_{6}(2510)$ $1689.52$ $31.5\%$ $1917.12$ $22.2\%$
$1^{+}(1^{+-})$ $b_{1}(1235)$ $1229.5$ $0\%$ $1229.5$ $0\%$
$1^{-}(2^{-+})$ $\pi_{2}(1670)$ $1527.41$ $8.57\%$ 239.41 $J=(1.03\pm 0.05)M^{2}-0.48\pm 0.15$ $1738.78$ $4.08\%$
$1^{+}(3^{+-})$ $b_{3}(2030)$ $1816.70$ $10.6\%$ $2129.56$ $4.80\%$
$1^{-}(4^{-+})$ $\pi_{4}(2250)$ $2099.95$ $6.67\%$ $2459.00$ $9.29\%$
$0^{-}(1^{+-})$ $h_{1}(1170)$ $1166.0$ $0\%$ $1166$ $0\%$
$0^{+}(2^{-+})$ $\eta_{2}(1645)$ $1448.56$ $10.4\%$ 227.04 $J=(1.15\pm 0.05)M^{2}-0.48\pm 0.15$ $1648.97$ $1.98\%$
$0^{-}(3^{+-})$ $h_{3}(2025)$ $1722.87$ $14.9\%$ $2019.57$ $0.27\%$
$0^{+}(4^{-+})$ $\eta_{4}(2330)$ $1991.49$ $14.5\%$ $2332.00$ $0.17\%$
Table 2: Masses in MeV for light unflavored mesons classified under even or odd $PC$, with isospins $I=0,1$, and spins $J=S\geq 1$ from the HW model, Eq. (16), and their corresponding Regge trajectories in GeV, and $\Lambda_{\rm HW}$ in MeV for each family. For comparison, we also show masses from the SW model, according to $m_{S}^{2}\propto S$ Karch:2006pv . The percentage deviations of both models refer to PDG data presented in Table 1.
@@
$J_{\nu}\bigg{(}\frac{m_{\nu,1}}{\Lambda_{\rm HW}}\bigg{)}=0.$ (15)
@@
$m_{\nu,1}=\chi_{\nu,1}\Lambda_{\rm HW};\qquad\qquad(\nu=S+1).$ (16)
@@
taking $m_{2,1}$ as an input from PDG PDG:2022pth for the states $\rho(770)$, $\omega(782)$, $b_{1}(1235)$, and $h_{1}(1170)$. The meson masses obtained from the HW model, using Eq. (16), are shown in Table 2, with the corresponding values of $\Lambda_{\rm HW}$ and Regge trajectories found for each mesonic family, together with the SW model predictions with the relative deviations with respect to PDG data, presented in Table 1.
Figure 1: Regge trajectories ($J\times M^{2}$) from the HW model (blue lines) and from PDG data (orange lines) for mesons, with masses expressed in GeV. Red dots are the values obtained from Eq. (16) in the HW model with uncertainties shown in Table 2, while the black dots correspond to PDG data shown in Table 1. The blue and orange lines are obtained by linear regression of these points. Upper left panel: mesons with even ${PC}$, $I=1$ ($\rho(770)$ family); Upper right panel: mesons with even ${PC}$, $I=0$ ($\omega(782)$ family); Lower left panel: mesons with odd ${PC}$, $I=1$ ($b_{1}(1235)$ family); Lower right panel: mesons with odd ${PC}$, $I=0$ ($h_{1}(1170)$ family).
@@
The process described above is known as the holographic HW model and was used in the past to obtain hadronic spectra among other properties Polchinski:2001tt ; Polchinski:2002jw ; Boschi-Filho:2002xih ; Boschi-Filho:2002wdj ; deTeramond:2005su ; Erlich:2005qh ; Boschi-Filho:2005xct ; Boschi-Filho:2005nmp ; Boschi-Filho:2006hfm ; Herzog:2006ra ; BallonBayona:2007vp ; Grigoryan:2007my ; Kwee:2007dd ; Grigoryan:2007wn ; Jo:2009xr ; Costa:2013uia ; Craps:2013iaa ; Mamo:2015dea ; Ballon-Bayona:2017bwk . Nevertheless, it is a well-known fact that this model implies non-linear Regge trajectories Boschi-Filho:2002xih ; Boschi-Filho:2002wdj ; deTeramond:2005su ; Erlich:2005qh ; Boschi-Filho:2005xct ; Karch:2006pv in contrast with experimental data (see, e.g., Spector:1967xsn ; Kanki:1975fp )
For comparison, we plot in Figure 1 experimental data (from Table 1) for the meson families starting with the states $\rho(770)$, $\omega(782)$, $b_{1}(1235)$, and $h_{1}(1170)$ together with HW masses (shown in Table 2) with uncertainties and the corresponding Regge trajectories obtained by linear regression of these points. The equations for these trajectories are also shown in Tables 1 and 2. As one can see from this Figure, the agreement between HW and PDG masses is not very good, specially for the $\rho(770)$ and $\omega(782)$ families.
@@
In the next section we develop an improvement of the HW model, by considering anomalous dimensions, in order to obtain better masses compared with the PDG ones, and Regge trajectories for these meson families which are closer to experimental data.